The following book and fair values were available for Westmont Company as of March 1. Book Value Fair Value Inventory $ 406,000 $ 363,500 Land 817,500 1,087,500 Buildings 2,005,000 2,314,750 Customer relationships 0 868,500 Accounts payable (128,500 ) (128,500 ) Common stock (2,000,000 ) Additional paid-in capital (500,000 ) Retained earnings, 1/1 (424,500 ) Revenues (482,500 ) Expenses 307,000

Arturo Company pays $4,150,000 cash and issues 22,800 shares of its $2 par value common stock (fair value of $50 per share) for all of Westmont’s common stock in a merger, after which Westmont will cease to exist as a separate entity. Stock issue costs amount to $33,500 and Arturo pays $44,000 for legal fees to complete the transaction.

Prepare Arturo’s journal entries to record its acquisition of Westmont. (If no entry is required for a transaction/event, select "No journal entry required" in the first account field.)
Answer #1

Answer:

No General Journal Debit Credit
1 Inventory           363,500
Land        1,087,500
Buildings        2,314,750
Customer relationships           868,500
Goodwill           784,250
Accounts payable      128,500
Common Stock        45,600
Cash 4,150,000
Additional Paid-In Capital 1,094,400
To record its acquisition of Westmont
2 Professional services expense 44,000
Cash 44,000
To record legal fees
3 Additional Paid-In Capital 33,500
       Cash 33,500
To record payment of stock issuance costs.

Calculation:

Entry #1:

To record its acquisition of Westmont:

Common Stock = 22800 x 2 = 45,600

Additional Paid-In Capital = 22800 x (50 - 2) = 1,094,400

Goodwill = Inventory + Land + Buildings + Customer relationships - Accounts payable - Common Stock - Cash - Additional Paid-In Capital = 363,500 + 1,087,500 + 2,314,750 + 868,500 - 128,500 - 45,600 - 4,150,000 - 1,094,400 = 784,250

The other two entries doesnt require any calculation since the amounts are provided in the question.